# Towards Fair and Trustworthy Agent-to-Agent Negotiations in Consumer Settings

### v1

@@ 4 Understanding Economic Risks for Real-World Users
Autonomous AI agents could potentially bring huge economic value to the users in many settings. However, they may also introduce systematic risks when being deployed at scale  (Feliu, 2001; Jabłonowska et al., 2018; Rohden and Zeferino, 2023; Deng et al., 2025; Hammond et al., 2025; Chen et al., 2025). In this section, we discuss the potential risks when both buyers and sellers delegate their negotiations and decision-making to AI agents and how models’ capability gaps and anomalies may translate into tangible economic losses for real users.
@@ 4.2 From Model Anomaly to Financial Risks
Fully automated, agent-based negotiation systems are prone to various anomalies stemming from unstable decision-making and imperfect instruction following of their base LLMs (Lan et al., 2025; Zhang et al., 2025; Cemri et al., 2025). While such failures may seem trivial or expected in research settings, they pose tangible risks to users in real-world settings. In this section, we analyze four model behavioral anomalies, pinpoint the conditions that trigger them, and outline how they can be translated into real financial loss for users. The detailed mathematical formula for the following anomaly measurement can be found in Appendix B.2.
@@
Constraint Violation.
Consider scenarios where a user authorizes an AI agent to negotiate on their behalf with a fixed budget $\beta$. If the agent accepts a deal above the budget, the agent may overdraw the account or exceed the user’s willingness to pay. Similarly, a seller agent agreeing to prices below the seller’s cost $p_{w}$ incurs guaranteed losses. In our evaluation, we quantify such anomaly using two metrics: Out-of-Budget Rate ($OBR$) and Out-of-Wholesale Rate ($OWR$). As shown in Figure 7, for $OBR$, we find that models with stronger negotiation capabilities, such as the DeepSeek series and Latest Generation GPT Series, including GPT-4.1, o4-mini, o3 and GPT-4o-mini, generally respect budget constraints and reject infeasible deals as buyers. However, models like GPT-3.5 and Qwen-7B frequently breach constraints, accepting deals above their budget in over 10% of all cases. This issue occurs across different budget settings but becomes even more serious when the user has a relatively low budget, posing serious risks for users in bad financial situations.
@@
Excessive payment.
Our experiments uncover a surprising anomaly: in some cases, buyer agents pay more than the listed retail price. We quantify this behavior with Overpayment Rate (OPR)–the proportion of successful deals in which the final transaction price is higher than the retail price despite the buyer’s budget affording a lower amount. As shown by Figure 8 (Left), overpayment often occurs under high-budget settings. Except for the DeepSeek family and Latest Generation GPT Series (GPT-4.1, o4-mini and o3), every model overpays to some degree when buyers have large $\beta$ values. To further investigate this issue, we qualitatively examined a sample of the negotiation history. As illustrated in Figure 8 (Right), overpayment often occurs after sellers ask buyers to reveal their budget early in the conversation. Despite our system prompt explicitly instructing buyers not to disclose their budget unless strictly necessary, many buyer agents reveal their budget easily. Sellers then try to anchor their offers to the disclosed number, even when it is higher than the listing price, and buyers accept the inflated deal without any objection.
Negotiation Deadlock.
Imagine a user who uses an API-based buyer agent for negotiation, expecting it to operate efficiently within reasonable bounds. A lay user would typically assume the agent will either reach a deal or end the negotiation in proper situations. However, in our experiment, we observe that agents might continue bargaining even when the seller has clearly stated a firm bottom line, leading to unnecessarily long negotiations. This behavior wastes computational resources and undermines the practical utility of automation. Here we name this issue as “Negotiation Deadlock” and formally define a negotiation deadlock as any dialogue that reaches the maximum number of rounds $T_{\text{max}}$ without a final agreement or explicit rejection.
@@
When analyzing buyer agents under different budget constraints, we observe a notable phenomenon that may cause the buyer to overpay: as the budget increases, particularly at or above the retail price, models tend to accept the seller’s proposed price as soon as it is below the budget rather than striving for better prices. In contrast, lower budgets (i.e., below retail price) appear to stimulate stronger bargaining behaviors, resulting in higher average price reduction rates $PRR_{\text{Buyer}}$. As shown in Figure 10, $PRR_{\text{Buyer}}$ exhibits a clear downward trend as the buyer budget increases, with a gap of nearly 9% between the highest and lowest price reduction rates. In practical deployments, buyer agents may derive their negotiation strategy from user-provided financial context, such as account balances or spending limits. If higher available funds systematically reduce the agent’s bargaining effort, users with generous budgets could end up consistently overpaying, not due to market necessity but because the agent passively accepts prices without seeking better deals.
5 Related work
5.1 AI Negotiations
@@
Negotiation plays a key role in modern society, and many researchers have studied negotiations from different perspectives  (Rubinstein, 1982; Nash, 2024; Hua et al., 2024; Mensfelt et al., 2024). Due to the potential business value and the rising capabilities of deep neural networks, researchers and practitioners have explored methods to build automated negotiation models (Zhou et al., 2016; Lewis et al., 2017; He et al., 2018; Bakker et al., 2019). More recently, large language models (LLMs) have shown strong capabilities in contextual understanding and strategic generation, leading to a growing interest in prompt-based LLM agents for complex negotiation tasks (Abdelnabi et al., 2024; Schneider et al., 2024; Bianchi et al., 2024; Shea et al., 2024; Yang et al., 2024b, 2025).
5.2 AI Agent in Consumer Settings
@@
A growing body of research examines AI agents in consumer-facing contexts, focusing on trust, decision delegation, and behavioral responses. Prior work has studied how agent intelligence and anthropomorphism shape consumer trust (Song and Lin, 2024; Zhao et al., 2025), and how task type affects willingness to delegate decisions (Frank et al., 2021; Fan and Liu, 2022; Yao et al., 2025). Chatbots and similar agents have also been explored as service intermediaries that influence consumer experience and perceived agency (Chong et al., 2021). While these studies offer important insights, they largely view agents as passive advisors or interfaces. Recent work begins to explore more active roles: ACE (Shea et al., 2024) introduces a negotiation training environment for LLM agents, and FishBargain (Kong et al., 2025) develops a seller-side bargaining agent for online flea markets. However, few studies systematically analyze how consumer-side agents negotiate with business agents, or how agent capabilities shape negotiation outcomes in real scenarios.
6 Discussion
@@
7 Conclusion
Along with the large-scale deployment of AI agents in real consumer settings, agent-to-agent interactions will become ubiquitous in the near future. But what will happen when we fully automate negotiation and deal-making with consumer and seller authorized AI agents? In this paper, we designed an experimental framework to investigate potential issues and risks in Agent-to-Agent negotiations and transactions. Our analysis reveals that Agent-to-Agent negotiation and transaction is naturally an imbalanced game where users using less capable agents will face significant financial loss against stronger agents. Furthermore, we found that LLMs’ anomaly might be transferred to real economic loss when they are deployed in real consumer settings. Our paper highlights the potential risks of using LLM agents to automate negotiation and transactions in real consumer settings.
Acknowledgement
We thank the generous funding support from the Project Liberty. We thank Yijia Shao, Jared Moore, Zach Robertson, Andreas Haupt, and Sophia Kazinnik for their kind feedback on the paper.
Appendix A Details of Dataset
A.1 Data Structure
@@
A.2 Wholesale Generation Prompt
@@
Appendix B Details of Metrics.
B.1 Main
Price Reduction Rate(PRR).
The Price Reduction Rate(PRR) quantifies the relative price change achieved through negotiation:
@@ B.1 Main
Relative Profit (RP).
We define the Relative Profit (RP) as the ratio between the total profit achieved by the model and the minimum reference profit (e.g. the GPT-3.5 profit in main experiment):
@@
Deal Rate(DR).
The Deal Rate (DR) measures the percentage of negotiations that result in a successful transaction:
@@
Profit Rate (PR).
We define the Profit Rate (PR) as the average per-product profit margin across all successful transactions. For each deal, the profit margin is computed relative to the wholesale cost. Formally:
@@
B.2 Anomaly
Out of Budget Rate (OBR).
The Out of Budget Rate (OBR) quantifies how often the final accepted price exceeds the buyer’s budget constraint:
@@ B.2 Anomaly
Out of Wholesale Rate (OWR).
The Out of Wholesale Rate (OWR) measures how often the final accepted price falls below the wholesale price, indicating unprofitable transactions from the seller’s perspective:
@@
Overpayment Rate (OPR).
The Overpayment Rate (OPR) quantifies how often the buyer ends up paying more than the reference retail price of the product in a successful transaction:
@@
Deadlock Rate (DLR).
The Deadlock Rate (DLR) quantifies the proportion of negotiations that reach the maximum allowed number of rounds $T_{\max}$ without reaching any agreement:
@@
Appendix C Details of Negotiation Implementation
C.1 System Prompt of Buyer
@@ C.2 Greeting Prompt
C.3 System Prompt of Seller
@@
C.4 System Prompt of Judge
@@
C.5 System Prompt of Analyst
@@
Appendix D Details of More Results
D.1 Understanding the Negotiation Gap via Model Specifications and Common Benchmarks.
@@ Appendix D Details of More Results
To investigate the sources of variation in negotiation capacity across models, we collect data on four commonly referenced model characteristics as potential explanatory factors. [^2] We obtain these data from model providers’ official websites or technical papers: https://openai.com/index/hello-gpt-4o/; https://arxiv.org/abs/2501.12948; https://qwenlm.github.io/blog/qwen2.5-llm/. The parameter count for GPT-4o-mini is estimated based on analysis in https://arxiv.org/abs/2412.19260. , including one architectural attribute: model size (in billions of parameters), and three performance-based benchmarks: general task performance (MMLU [Hendrycks et al., 2020]), mathematical ability (MATH [Hendrycks et al., 2021]), and scientific ability (GPQA [Rein et al., 2024]). We combine three negotiation-relevant metrics—Buyer Price Reduction Rate($PRR_{\text{Buyer}}$), reverse of Seller Price Reduction($1-PRR_{\text{Seller}}$), and $RP$—into a scalar indicator via z-score normalization followed by averaging, yielding a composite Negotiation Capacity Score ($NCS$). We then compute the Pearson correlation between each model’s $NCS$ and the four benchmark scores. As shown in Figure 19, negotiation capacity shows a very strong correlation with general task performance on MMLU ($r=0.93$), along with substantial correlations with mathematical ($r=0.87$) and scientific ability ($r=0.80$). The weakest correlation appears with model size ($r=0.53$), which we attribute to multiple factors: some high-parameter models (e.g., GPT-3.5) belong to earlier generations with less optimized architectures and performance, while for commercial models such as GPT-4o-mini, exact parameter counts are unavailable and must be estimated from external sources.
D.2 Negotiation Capacity Gap Indicates Behavioral Robustness Gap.
@@
Figures 7, 8, and 9 present anomaly indicators across six models analyzed in Section 3.2. The data reveals a notable pattern: the proportion of anomalies appears inversely related to the models’ negotiation capabilities. This observation motivates the research question: Are models with stronger negotiation skills also more robust against automation-induced anomalies?
To investigate this relationship, we reuse the previously defined Negotiation Capacity Score (NCS) (see Section D.1). To quantify a model’s overall tendency toward negotiation anomalies, we construct a composite Risk Index by aggregating the four anomaly-related indicators introduced in Section 4.2. Each indicator is standardized using z-score normalization and averaged to produce a unified scalar value. We then compute the Pearson correlation between NCS and the Risk Index. As shown in Figure 20, the result ($r=-0.67$) indicates a moderate negative association: models with higher negotiation capacity consistently exhibit lower anomaly indices, suggesting greater behavioral robustness in automated negotiation systems.

### v3

@@ 4 Understanding Economic Risks for Real-World Users
Autonomous AI agents could potentially bring huge economic value to the users in many settings. However, they may also introduce systematic risks when being deployed at scale  (feliu2001intelligent; jablonowska2018consumer; rohden2023recommendation; deng2025ai; hammond2025multi; chen2025position). In this section, we discuss the potential risks when both buyers and sellers delegate their negotiations and decision-making to AI agents and how models’ capability gaps and anomalies may translate into tangible economic losses for real users.
@@ 4.2 From Model Anomaly to Financial Risks
Fully automated, agent-based negotiation systems are prone to various anomalies stemming from unstable decision-making and imperfect instruction following of their base LLMs (lan2025exploring; zhang2025agent; cemri2025multi). While such failures may seem trivial or expected in research settings, they pose tangible risks to users in real-world settings. In this section, we analyze four model behavioral anomalies, pinpoint the conditions that trigger them, and outline how they can be translated into real financial loss for users. The detailed mathematical formula for the following anomaly measurement can be found in Appendix E.2.
@@
Constraint Violation. Consider scenarios where a user authorizes an AI agent to negotiate on their behalf with a fixed budget $\beta$. If the agent accepts a deal above the budget, the agent may overdraw the account or exceed the user’s willingness to pay. Similarly, a seller agent agreeing to prices below the seller’s cost $p_{w}$ incurs guaranteed losses. In our evaluation, we quantify such anomaly using two metrics: Out-of-Budget Rate ($OBR$) and Out-of-Wholesale Rate ($OWR$). As shown in Figure 6, for $OBR$, we find that models with stronger negotiation capabilities, such as the DeepSeek series and Latest Generation GPT Series, including GPT-4.1, o4-mini, o3 and GPT-4o-mini, generally respect budget constraints and reject infeasible deals as buyers. However, models like GPT-3.5 and Qwen-7B frequently breach constraints, accepting deals above their budget in over 10% of all cases. This issue occurs across different budget settings but becomes even more serious when the user has a relatively low budget, posing serious risks for users in bad financial situations.
@@
Excessive payment. Our experiments uncover a surprising anomaly: in some cases, buyer agents pay more than the listed retail price. We quantify this behavior with Overpayment Rate (OPR)–the proportion of successful deals in which the final transaction price is higher than the retail price despite the buyer’s budget affording a lower amount. As shown by Figure 7 (Left), overpayment often occurs under high-budget settings. Except for the DeepSeek family and Latest Generation GPT Series (GPT-4.1, o4-mini and o3), every model overpays to some degree when buyers have large $\beta$ values. To further investigate this issue, we qualitatively examined a sample of the negotiation history. As illustrated in Figure 7 (Right), overpayment often occurs after sellers ask buyers to reveal their budget early in the conversation. Despite our system prompt explicitly instructing buyers not to disclose their budget unless strictly necessary, many buyer agents reveal their budget easily. Sellers then try to anchor their offers to the disclosed number, even when it is higher than the listing price, and buyers accept the inflated deal without any objection.
Negotiation Deadlock. Imagine a user who uses an API-based buyer agent for negotiation, expecting it to operate efficiently within reasonable bounds. A lay user would typically assume the agent will either reach a deal or end the negotiation in proper situations. However, in our experiment, we observe that agents might continue bargaining even when the seller has clearly stated a firm bottom line, leading to unnecessarily long negotiations. This behavior wastes computational resources and undermines the practical utility of automation. Here we name this issue as “Negotiation Deadlock” and formally define a negotiation deadlock as any dialogue that reaches the maximum number of rounds $T_{\text{max}}$ without a final agreement or explicit rejection.
@@
Early Settlement. When analyzing buyer agents under different budget constraints, we observe a notable phenomenon that may cause the buyer to overpay: as the budget increases, particularly at or above the retail price, models tend to accept the seller’s proposed price as soon as it is below the budget rather than striving for better prices. In contrast, lower budgets (i.e., below retail price) appear to stimulate stronger bargaining behaviors, resulting in higher average price reduction rates $PRR_{\text{Buyer}}$. As shown in Figure 9, $PRR_{\text{Buyer}}$ exhibits a clear downward trend as the buyer budget increases, with a gap of nearly 9% between the highest and lowest price reduction rates. In practical deployments, buyer agents may derive their negotiation strategy from user-provided financial context, such as account balances or spending limits. If higher available funds systematically reduce the agent’s bargaining effort, users with generous budgets could end up consistently overpaying, not due to market necessity but because the agent passively accepts prices without seeking better deals.
5 Conclusion
With the growing deployment of AI agents in consumer settings, automated agent-to-agent negotiations will become widespread. We present an experimental framework to explore potential risks in such interactions. Our analysis shows that these negotiations inherently favor stronger agents, causing financial losses for users with weaker agents. Additionally, anomalies in LLM behavior can translate into real economic harm. Our study underscores significant risks associated with automating consumer negotiations through LLM agents.
Appendix A Appendix
Appendix B Related work
B.1 AI Negotiations
@@
Negotiation plays a key role in modern society, and many researchers have studied negotiations from different perspectives  (rubinstein1982perfect; nash2024non; hua2024game; mensfelt2024autoformalizing). Due to the potential business value and the rising capabilities of deep neural networks, researchers and practitioners have explored methods to build automated negotiation models (zhou2016multiagent; lewis2017deal; he2018decoupling; bakker2019rlboa). More recently, large language models (LLMs) have shown strong capabilities in contextual understanding and strategic generation, leading to a growing interest in prompt-based LLM agents for complex negotiation tasks (abdelnabi2024cooperation; schneider2024negotiating; bianchi2024well; shea2024ace; yang2024makes; yang2025fraud).
B.2 AI Agent in Consumer Settings
@@
A growing body of research examines AI agents in consumer-facing contexts, focusing on trust, decision delegation, and behavioral responses. Prior work has studied how agent intelligence and anthropomorphism shape consumer trust (song2024exploring; zhao2025consumers), and how task type affects willingness to delegate decisions (frank2021ai; fan2022exploring; yao2025your). Chatbots and similar agents have also been explored as service intermediaries that influence consumer experience and perceived agency (chong2021ai). While these studies offer important insights, they largely view agents as passive advisors or interfaces. Recent work begins to explore more active roles: ACE (shea2024ace) introduces a negotiation training environment for LLM agents, and FishBargain (kong2025fishbargain) develops a seller-side bargaining agent for online flea markets. However, few studies systematically analyze how consumer-side agents negotiate with business agents, or how agent capabilities shape negotiation outcomes in real scenarios.
Appendix C Discussion
@@
Appendix D Details of Dataset
D.1 Data Structure
@@
D.2 Wholesale Generation Prompt
@@
Appendix E Details of Metrics.
E.1 Main
Price Reduction Rate(PRR). The Price Reduction Rate(PRR) quantifies the relative price change achieved through negotiation:
@@ E.1 Main
Relative Profit (RP). We define the Relative Profit (RP) as the ratio between the total profit achieved by the model and the minimum reference profit (e.g. the GPT-3.5 profit in main experiment):
@@
Deal Rate(DR). The Deal Rate (DR) measures the percentage of negotiations that result in a successful transaction:
@@
Profit Rate (PR). We define the Profit Rate (PR) as the average per-product profit margin across all successful transactions. For each deal, the profit margin is computed relative to the wholesale cost. Formally:
@@
E.2 Anomaly
Out of Budget Rate (OBR). The Out of Budget Rate (OBR) quantifies how often the final accepted price exceeds the buyer’s budget constraint:
@@ E.2 Anomaly
Out of Wholesale Rate (OWR). The Out of Wholesale Rate (OWR) measures how often the final accepted price falls below the wholesale price, indicating unprofitable transactions from the seller’s perspective:
@@
Overpayment Rate (OPR). The Overpayment Rate (OPR) quantifies how often the buyer ends up paying more than the reference retail price of the product in a successful transaction:
@@
Deadlock Rate (DLR). The Deadlock Rate (DLR) quantifies the proportion of negotiations that reach the maximum allowed number of rounds $T_{\max}$ without reaching any agreement:
@@
Appendix F Details of Negotiation Implementation
F.1 System Prompt of Buyer
@@ F.2 Greeting Prompt
F.3 System Prompt of Seller
@@
F.4 System Prompt of Judge
@@
F.5 System Prompt of Analyst
@@
Appendix G Details of More Results
G.1 Understanding the Negotiation Gap via Model Specifications and Common Benchmarks.
@@ Appendix G Details of More Results
To investigate the sources of variation in negotiation capacity across models, we collect data on four commonly referenced model characteristics as potential explanatory factors. [^2] We obtain these data from model providers’ official websites or technical papers: https://openai.com/index/hello-gpt-4o/; https://arxiv.org/abs/2501.12948; https://qwenlm.github.io/blog/qwen2.5-llm/. The parameter count for GPT-4o-mini is estimated based on analysis in https://arxiv.org/abs/2412.19260. , including one architectural attribute: model size (in billions of parameters), and three performance-based benchmarks: general task performance (MMLU (hendrycks2020measuring)), mathematical ability (MATH (hendrycksmath2021)), and scientific ability (GPQA (rein2024gpqa)). We combine three negotiation-relevant metrics—Buyer Price Reduction Rate($PRR_{\text{Buyer}}$), reverse of Seller Price Reduction($1-PRR_{\text{Seller}}$), and $RP$—into a scalar indicator via z-score normalization followed by averaging, yielding a composite Negotiation Capacity Score ($NCS$). We then compute the Pearson correlation between each model’s $NCS$ and the four benchmark scores. As shown in Figure 18, negotiation capacity shows a very strong correlation with general task performance on MMLU ($r=0.93$), along with substantial correlations with mathematical ($r=0.87$) and scientific ability ($r=0.80$). The weakest correlation appears with model size ($r=0.53$), which we attribute to multiple factors: some high-parameter models (e.g., GPT-3.5) belong to earlier generations with less optimized architectures and performance, while for commercial models such as GPT-4o-mini, exact parameter counts are unavailable and must be estimated from external sources.
G.2 Negotiation Capacity Gap Indicates Behavioral Robustness Gap.
@@
Figures 6, 7, and 8 present anomaly indicators across six models analyzed in Section 3.2. The data reveals a notable pattern: the proportion of anomalies appears inversely related to the models’ negotiation capabilities. This observation motivates the research question: Are models with stronger negotiation skills also more robust against automation-induced anomalies?
To investigate this relationship, we reuse the previously defined Negotiation Capacity Score (NCS) (see Section G.1). To quantify a model’s overall tendency toward negotiation anomalies, we construct a composite Risk Index by aggregating the four anomaly-related indicators introduced in Section 4.2. Each indicator is standardized using z-score normalization and averaged to produce a unified scalar value. We then compute the Pearson correlation between NCS and the Risk Index. As shown in Figure 19, the result ($r=-0.67$) indicates a moderate negative association: models with higher negotiation capacity consistently exhibit lower anomaly indices, suggesting greater behavioral robustness in automated negotiation systems.